# Mass-imbalance effect on the cluster formation in a one-dimensional Fermi gas with coexistent $s$- and $p$-wave interactions

### v1

@@ Abstract
We consider the mass-imbalance effect on the clustering in a one-dimensional two-component Fermi gas with coexistent even- and odd-wave interactions resulting in different configurations of clustering phases. We obtain the solutions of both stable two- and three-body cluster states with different mass ratios and configurations by solving the corresponding variational equations. We feature out phase diagrams consisting of the $s$- and $p$-wave pairing phases, and tripling phase with different configurations, in a plane of $s$- and $p$-wave pairing strengths. As for the in-vacuum case, the three-body clustering is always the lowest-lying phase. While for the in-medium case, the Cooper tripling phase dominates over the pairing phases when both $s$- and $p$-wave interactions are moderately strong. There is also a competition between different clustering configurations of three-body clustering.
@@ I Introduction
It would be useful to study the mass-imbalance effect on multibody clustering in various systems. The Efimov state [6] has attracted lots of interest in ultracold atoms, nuclear physics, and so on. In the corresponding investigation, experimental evidence of Efimov physics has been reported in different mass-imbalanced ultracold atomic gas mixtures such as 41K-87Rb [7] or 6Li-133Cs [8] ones. The two-nucleon halo nuclei such as 11Li (9Li core ++ 22 halo neutrons) is also regarded as an ideal candidate to experimentally detect the Efimov states [9]. So far, various possible superconductors with competing odd- and even-parity pairing channels have been experimentally realized in the mass-imbalanced systems, including UTe2 [4, 10], UPt3 [11], WTe2 [13], and Sr2RuO4 [14, 15]. Moreover, in nuclear physics, there is also a novel state called hypernucleus consisting of at least one hyperon in addition to the normal nucleons [16]. It shows many interesting properties, such as that hypernuclei containing the lightest hyperon $\Lambda$ (such as $^{7}_{\Lambda}$Li) tend to be more tightly bound than normal nuclei [17]. The hypernuclei with small particle number would be useful for the preliminary studies.
@@
The ultracold atomic gas is regarded as an ideal platform for investigating such non-trivial quantum many-body phenomena, due to its tunnable interactions and densities [18, 19]. Both $s$- and $p$-wave interactions are tunable near the Feshbach resonance [20, 21, 22, 23, 19, 24, 25]. Consequently, it is interesting to realize the situation that Fermi superfluids with hybridized $s$- and $p$-wave pairings in cold atomic systems via overlapped Feshbach resonances [18, 26, 27]. In such a system, in addition to $s$- and $p$-wave Cooper pairings, the clustering associated with the Cooper instability may be extended to induce the bound states consisting of three particles or even more. In order to study such a many-body problem, the generalized Cooper problem has been further applied to cluster states such as three-body Cooper triples [28, 29, 30, 31, 32, 5, 33] and even four-body Cooper quartets [34, 35, 36, 37, 38, 39]. On the other hand, with the help of specific geometry and trap potential, low dimensions and lattice structures can be achieved in cold-atomic systems [40, 41, 42, 43]. Such low-dimensional systems in condensed-matter or nuclear physics have also attracted much attention [44, 45, 46].
@@
In the previous work [5], we studied the degenerate two-component fermions with coexistent interspecies and intraspecies interactions, which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the Cooper triple with mass-balanced configuration. However, the mass imbalance is also worth being investigated, and the intraspecies interaction is not necessary limited to exist only within one component. In this work, we are going to figure out the properties in a system with both $s$- and $p$-wave interactions as shown in Fig. 1. The two- and three-body clusterings with different configurations are also anticipated. In Ref. [47], by activating the orbitals degree of freedom, the authors reported that the emergence of even-parity ($s$-wave) interaction in a spin-polarized fermionic 40K gas besides the odd-wave ($p$-wave) interactions among identical fermions in each orbital. Based on that, present model setup might be achieved experimentally via fulfilling multiorbital structure with different components of atoms in the future.
@@
Figure 1: Schematic figure representing our model. We consider the degenerate two-component mass-imbalanced fermions (states $a$ and $b$ occupying the levels with momentum $k$ up to the Fermi momentum $k_{\rm F}$) and the consequence of coexistent interspecies $s$-wave interaction $V_{s}$ and intraspecies $p$-wave interactions $V_{i}$ (acting on two identical fermions in the $i=a,b$ state), which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the Cooper triples.
@@
In this paper, we investigate the fate of in-vacuum and in-medium two- and three-body correlations in a one-dimensional two-component Fermi gas with coexistent $s$- and $p$-wave interactions. Specific attention is paid to the mass imbalance and different configurations of clustering. In general, the mass of two components of fermions are adopted as different ones, and $p$-wave interactions among both kinds of atoms are taken into considerations. By solving the variational equations, we show the solutions of the stable clusters. We also figure out a phase diagram consisting of pairing and tripling states with different configurations.
@@
This paper is organized as follows. The theoretical framework is presented in Sec. II, where we show the Hamiltonian for present system. We apply a variational approach for in-medium two- and three-body states on top of the Fermi sea to this model. In Sec. III.1, we first discuss our numerical results for the in-vacuum bound states and the ground-state phase diagram. The results of in-medium case are shown in Sec. III.2. Finally, a summary and perspectives will be given in Sec. IV. In the following, we take $\hbar=c=k_{\rm B}=1$. The system size is taken to be a unit.
@@ II Theoretical framework
In this work, we consider a one-dimensional mass-imbalanced two-component fermions with coexistent short-range interactions. The intraspecies odd-parity ($p$-wave) interactions among each kind of fermions are both taken into considerations. While as for the interspecies interaction, we assume the even-parity ($s$-wave) one is dominant and neglect the odd-parity ($p$-wave) interaction one as the same as the previous work [5]. Consequently, the present system can be described by the following Hamiltonian as
@@
where $\xi_{k,i}=k^{2}/(2m_{i})-\mu_{i}$ ($i=a,b$) in the kinetic term $K$ is the single-particle energy with a momentum $k$, an atomic mass $m_{i}$, and chemical potential $\mu_{i}$. Here the reduced mass reads $m_{r}=\frac{m_{a}m_{b}}{m_{a}+m_{b}}$. Note that in general the cases with $m_{a}\neq m_{b}$ will be considered in this paper. In the generalized Cooper problems, we take $\mu_{i}=E_{\textrm{F},i}$ where $E_{\textrm{F},i}$ is the Fermi energy of species $i$. Due to the mass difference, there is a mismatch of the Fermi energy (or chemical potential) between two different kinds of fermions, namely, $\mu_{a}\neq\mu_{b}$. $V_{i}$ represents the two-body $p$-wave interaction between two identical atoms $i$ with a coupling constant $U_{i}$, and $V_{s}$ corresponds to the two-body $s$-wave interaction with a coupling constant $U_{s}$. The contact couplings $U_{s}$ and $U_{i}$ can be renormalized by introducing the $s$-wave and $p$-wave scattering lengths as
@@
As schematically shown in Fig. 1, with the mass imbalance and the $p$-wave interactions among both kinds of atoms, there are in total five phases of correlations expected to appear in the present system. In detail, as for the pairing states, we have $s$-wave pairing, $p$-wave pairing ($aa$ configuration) and $p$-wave pairing ($bb$ configuration). While for the tripling states, we have aabaab and abbabb configurations, respectively.
@@
for tripling (aabaab configuration) state [5], and
@@
for tripling (abbabb configuration) state, respectively. In the above trial wave functions, $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ denotes the variational parameter for the state with $mn\cdots$ configuration and a set of momenta $q_{1},q_{2},\cdots$, and $|{\rm FS}\rangle$ represents the Fermi sea. From the parity and symmetry, it is easy to find $\Omega^{iij}_{p,q}=-\Omega^{iij}_{-p,q}$ and $\Omega_{p,q}^{iij}=\Omega_{p,-q}^{iij}$ ($i,j=a,b$ and $i\neq j$) for Cooper triple phases. In addition, here $\sum_{p_{1},p_{2},\cdots}^{\prime}$ is adopted to denote the the momentum summation restricted by the Fermi surface for convenience. The step functions associated with the Fermi surface will be recovered when we numerically evaluate the momentum summation. By minimizing the in-medium two- and three-body ground-state energies based on the variational principle, the variational parameter $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ will be determined correspondingly.
@@
where $E_{\alpha}^{mn\cdots}$ ($\alpha=2,3$) is the ground-state energy of a Cooper cluster with $mn\cdots$ configuration. As a result,the in-medium variational equations read
@@
for the $s$-wave pairing [48],
@@
for tripling (aabaab configuration) [5], and
@@
for tripling (abbabb configuration), respectively. To simplify the ensuing expressions, in the above variational equations we introduce
@@
Before discussing the numerical demonstrations in detail, we note that present variational approach is able to describe the BCS-BEC crossover and its three-body counterpart qualitatively [49]. The variational wave functions for calculating the energies of Cooper clusters can well depict both the Cooper clustering in the weak-coupling BCS limit and the tightly bound state formation in the strong-coupling BEC limit in a unified manner.
@@ III Results and discussion
III.1 In-vaccum case
@@
Figure 2: (a). In-vacuum and in-medium $s$-wave pairing energies as functions of $s$-wave scattering length. (b). In-vacuum and in-medium $p$-wave pairing energies with different configurations as functions of $p$-wave scattering length. (c). In-vacuum and in-medium three-body energies with different configurations as functions of $p$-wave scattering length at fixed $s$-wave interaction strength as $1/(\Lambda a_{s})\times 10=2.7$ with $\Lambda$ the momentum cutoff. In all panels, mass ratio is fixed as $m_{a}/m_{b}=2$, and the reference energy scale is $E_{\Lambda}=\Lambda^{2}/(2m_{r})$.
@@
Figure 3: In-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$. The mass ratio is taken as $m_{a}=2m_{b}$.
We feature out the in-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$ as shown in Fig. 3. The in-vacuum phase diagram gives the lowest-lying phase among all the five configurations of two- or three-body clusters discussed in this work. The mass ratio is taken as $m_{a}=2m_{b}$. It is seen that in the interaction regime concerned in this paper, $|E_{3}|>|E_{2}|$ always holds regardless of specific two- and three-body configurations in the in-vacuum case. Moreover, at a fixed $s$-wave interaction, with the increase of $p$-wave interaction, the lowest-lying phase changes from the three-body states with aabaab configuration to those with abbabb configuration. In other words, the in-vacuum three-body states with abbabb configuration become dominant in the regime with stronger $p$-wave interaction. It is because the increase of $p$-wave attraction induces larger enhancement for $bb$ pairings with smaller mass compared to $aa$ ones.

### v2

@@ Abstract
We consider the mass-imbalance effect on the clustering in a one-dimensional two-component Fermi gas with coexistent even- and odd-wave interactions resulting in different configurations of clustering phases. We obtain the solutions of both stable two- and three-body cluster states with different mass ratios and configurations by solving the corresponding variational equations. We numerically map out phase diagrams consisting of the $s$- and $p$-wave pairing phases, and trimer phase with different configurations, in a plane of $s$- and $p$-wave pairing strengths. Within the explored ranges of $s$- and $p$-wave pairing strengths, the in-vacuum three-body states are always more deeply bound than the two-body ones. While for the in-medium case, the Cooper trimer phase dominates over the pairing phases when both $s$- and $p$-wave interactions are moderately strong. There is also a competition between different clustering configurations of three-body clustering.
@@ I Introduction
It would be useful to study the mass-imbalance effect on multibody clustering in various systems. The Efimov state [6] has attracted lots of interest in ultracold atoms, nuclear physics, and so on. In the corresponding investigation, experimental evidence of Efimov physics has been reported in different mass-imbalanced ultracold atomic gas mixtures such as 41K-87Rb [7] or 6Li-133Cs [8] ones. The two-nucleon halo nuclei such as 11Li (9Li core ++ two halo neutrons) is also regarded as an ideal candidate to experimentally detect the Efimov states [9]. So far, various possible superconductors with competing odd- and even-parity pairing channels have been experimentally realized in the mass-imbalanced systems, including UTe2 [4, 10], UPt3 [11], WTe2 [13], and Sr2RuO4 [14, 15]. Moreover, in nuclear physics, there is also a novel state called hypernucleus consisting of at least one hyperon in addition to the normal nucleons [16]. It shows many interesting properties, such as that hypernuclei containing the lightest hyperon $\Lambda$ (such as $^{7}_{\Lambda}$Li) tend to be more tightly bound than normal nuclei [17]. The hypernuclei with small particle number would be useful for the preliminary studies.
@@
The ultracold atomic gas is regarded as an ideal platform for investigating such non-trivial quantum many-body phenomena, due to its tunable interactions and densities [18, 19]. Both $s$- and $p$-wave interactions are tunable near the Feshbach resonance [20, 21, 22, 23, 19, 24, 25]. Consequently, it is interesting to realize the situation that Fermi superfluids with hybridized $s$- and $p$-wave pairings in cold atomic systems via overlapped Feshbach resonances [18, 26, 27]. In such a system, in addition to $s$- and $p$-wave Cooper pairings, the clustering associated with the Cooper instability may be extended to induce the bound states consisting of three particles or even more. In order to study such a many-body problem, the generalized Cooper problem has been further applied to cluster states such as three-body [28, 29, 30, 31, 32, 5, 33] and even four-body Cooper bound states [34, 35, 36, 37, 38, 39]. On the other hand, with the help of specific geometry and trap potential, low dimensions and lattice structures can be achieved in cold-atomic systems [40, 41, 42, 43]. Such low-dimensional systems in condensed-matter or nuclear physics have also attracted much attention [44, 45, 46].
@@
In the previous work [5], we studied the degenerate two-component fermions with coexistent interspecies and intraspecies interactions, which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the three-body Cooper bound states with mass-balanced configuration. In the following, we refer to the formation of three-body Cooper bound states as Cooper trimer formation. However, the mass imbalance is also worth being investigated, and the intraspecies interaction is not necessarily limited to exist only within one component. In this work, we are going to figure out the properties in a system with both $s$- and $p$-wave interactions as shown in Fig. 1. The two- and three-body clusterings with different configurations are also anticipated. In Ref. [47], by activating the orbitals degree of freedom, the authors reported that the emergence of even-parity ($s$-wave) interaction in a spin-polarized fermionic 40K gas besides the odd-wave ($p$-wave) interactions among identical fermions in each orbital. Based on that, present model setup might be achieved experimentally via fulfilling multiorbital structure with different components of atoms in the future.
@@
Figure 1: Schematic figure representing our model. We consider the degenerate two-component mass-imbalanced fermions (states $a$ and $b$ occupying the levels with momentum $k$ up to the Fermi momentum $k_{\rm F}$) and the consequence of coexistent interspecies $s$-wave interaction $V_{s}$ and intraspecies $p$-wave interactions $V_{i}$ (acting on two identical fermions in the $i=a,b$ state), which lead to the Cooper instabilities towards the $s$- and $p$-wave Cooper pairs, and the Cooper trimers.
@@
In this paper, we investigate the fate of in-vacuum and in-medium two- and three-body correlations in a one-dimensional two-component Fermi gas with coexistent $s$- and $p$-wave interactions. Specific attention is paid to the mass imbalance and different configurations of clustering. In general, the mass of two components of fermions are adopted as different ones, and $p$-wave interactions among both kinds of atoms are taken into account. By solving the variational equations, we show the solutions of the stable clusters. We also figure out a phase diagram consisting of pairing and trimer states with different configurations.
@@
This paper is organized as follows. The theoretical framework is presented in Sec. II, where we show the Hamiltonian for present system. We apply a variational approach for in-medium two- and three-body states on top of the Fermi sea to this model. In Sec. III.1, we first discuss our numerical results for the in-vacuum bound states and the lowest-energy cluster-instability diagram. The results of in-medium case are shown in Sec. III.2. Finally, a summary and perspectives will be given in Sec. IV. In the following, we take $\hbar=c=k_{\rm B}=1$. The system size is taken to be a unit.
@@ II Theoretical framework
In this work, we consider a one-dimensional mass-imbalanced two-component fermions with coexistent short-range interactions. The intraspecies odd-parity ($p$-wave) interactions among each kind of fermion are both taken into account. As for the interspecies interaction, we assume the even-parity ($s$-wave) one is dominant and neglect the odd-parity ($p$-wave) interaction one as the same as the previous work [5]. We note that in a two-component system this assumption does not restrict the possible three-body clustering configurations, and an interspecies odd-parity interaction would only introduce higher-order contributions within the same configurations. Consequently, the present system can be described by the following Hamiltonian as
@@
where $\xi_{k,i}=k^{2}/(2m_{i})-\mu_{i}$ ($i=a,b$) in the kinetic term $K$ is the single-particle energy with a momentum $k$, an atomic mass $m_{i}$, and chemical potential $\mu_{i}$. Here the reduced mass reads $m_{r}=\frac{m_{a}m_{b}}{m_{a}+m_{b}}$. Note that in general the cases with $m_{a}\neq m_{b}$ will be considered in this paper. In the generalized Cooper problems, we take $\mu_{i}=E_{\textrm{F},i}$ where $E_{\textrm{F},i}$ is the Fermi energy of species $i$. Namely, throughout the parameter scans, the chemical potentials are fixed by the Fermi energies $\mu_{i}=k_{\rm{F}}^{2}/(2m_{i})$, corresponding to equal densities but different masses. Due to the mass difference, there is a mismatch of the Fermi energy (or chemical potential) between two different kinds of fermions, namely, $\mu_{a}\neq\mu_{b}$. $V_{i}$ represents the two-body $p$-wave interaction between two identical atoms $i$ with a coupling constant $U_{i}$, and $V_{s}$ corresponds to the two-body $s$-wave interaction with a coupling constant $U_{s}$. The contact couplings $U_{s}$ and $U_{i}$ can be renormalized by introducing the $s$-wave and $p$-wave scattering lengths as
@@
As schematically shown in Fig. 1, with the mass imbalance and the $p$-wave interactions among both kinds of atoms, there are in total five phases of correlations expected to appear in the present system. In detail, as for the pairing states, we have $s$-wave pairing, $p$-wave pairing ($aa$ configuration) and $p$-wave pairing ($bb$ configuration). While for the trimer states, we have aabaab and abbabb configurations, respectively.
@@
for Cooper trimer (aabaab configuration) state [5], and
@@
for Cooper trimer (abbabb configuration) state, respectively. In the above trial wave functions, $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ denotes the variational parameter for the state with $mn\cdots$ configuration and a set of momenta $q_{1},q_{2},\cdots$, and $|{\rm FS}\rangle$ represents the Fermi sea. From the parity and symmetry, it is easy to find $\Omega^{iij}_{p,q}=-\Omega^{iij}_{-p,q}$ and $\Omega_{p,q}^{iij}=\Omega_{p,-q}^{iij}$ ($i,j=a,b$ and $i\neq j$) for Cooper trimer phases. In addition, here $\sum_{p_{1},p_{2},\cdots}^{\prime}$ is adopted to denote the momentum summation restricted by the Fermi surface for convenience. The step functions associated with the Fermi surface will be recovered when we numerically evaluate the momentum summation. By minimizing the in-medium two- and three-body energies based on the variational principle, the variational parameter $\Omega^{mn\cdots}_{q_{1},q_{2},\cdots}$ will be determined correspondingly. In the present work, we assume equal particle densities for the two components, such that $k_{{\rm F},a}=k_{{\rm F},b}\equiv k_{\rm F}$. As a consequence, Pauli blocking is implemented using a common Fermi momentum in Eqs. (8)–(II), while the chemical potentials satisfy $\mu_{i}=k_{\rm F}^{2}/(2m_{i})$ and differ due to the mass imbalance.
@@
where $E_{\alpha}^{mn\cdots}$ ($\alpha=2,3$) is the lowest-lying energy of a Cooper cluster with $mn\cdots$ configuration. As a result,the in-medium variational equations read
@@
for the $s$-wave pairing [Ohashi2020Prog.Part.andNucl.Phys.111.103739],
@@
for Cooper trimer (aabaab configuration) [5], and
@@
for Cooper trimer (abbabb configuration), respectively. To simplify the ensuing expressions, in the above variational equations we introduce
@@
Before discussing the numerical demonstrations in detail, we note that present variational approach is able to describe the BCS-BEC crossover and its three-body counterpart qualitatively [48]. The variational wave functions for calculating the energies of Cooper clusters can well depict both the Cooper clustering in the weak-coupling BCS limit and the tightly bound state formation in the strong-coupling BEC limit in a unified manner. We also emphasize that the present approach is a variational analysis for the generalized Cooper problem performed on top of an inert Fermi sea. Therefore, the obtained energies and phase diagrams should be understood as a hierarchy of lowest-energy cluster instabilities within the adopted variational subspace, rather than the fully self-consistent thermodynamic many-body ground state.
@@ III Results and discussion
In this section, we are going to present the numerical implementation of analytic derivations from Eqs. (19)–(27). All momentum summations are evaluated on uniform grids up to a finite cutoff $\Lambda$, with a grid spacing $\mathrm{d}k=0.01\,k_{\rm F}$, which is sufficient for numerical convergence. The coupled equations for the auxiliary amplitudes $\Gamma^{iij}_{p}(q)$ and $\Gamma^{iij}_{s}(k)$ are solved after discretization as a linear eigenvalue problem, whose lowest eigenvalue determines the cluster energy. The symmetry relations stated in Eq. (24) are explicitly enforced in the numerical implementation to reduce the independent degrees of freedom and improve numerical stability. Convergence is checked by refining the momentum grids, and by independently varying the cutoff $\Lambda$ while keeping the physical scattering parameters fixed.
III.1 In-vacuum case
@@
Figure 2: (a) In-vacuum and in-medium $s$-wave pairing energies as functions of the $s$-wave scattering length. (b) In-vacuum and in-medium $p$-wave pairing energies with different configurations as functions of the $p$-wave scattering length. (c) In-vacuum and in-medium three-body energies with different configurations as functions of the $p$-wave scattering length at a fixed $s$-wave interaction strength $1/(\Lambda a_{s})\times 10=2.7$. For the in-vacuum case, all energies are measured relative to the vacuum threshold. For the in-medium case, energies are measured relative to the energy of the filled Fermi sea of noninteracting fermions. A more strongly bound state corresponds to a more negative energy (or equivalently a larger $|E|$). In all panels, the mass ratio is fixed to $m_{a}/m_{b}=2$, and the momentum cutoff is $\Lambda/k_{F}=10$. The reference energy scale is $E_{\Lambda}=\Lambda^{2}/(2m_{r})$.
@@
Figure 3: In-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$. The phase shown corresponds to the most deeply bound cluster, determined by comparing the in-vacuum two- and three-body energies. All energies are measured relative to the vacuum threshold, and a more strongly bound state corresponds to a more negative energy. Within the explored parameter window, the three-body states are always more deeply bound than the two-body ones. The mass ratio is fixed to $m_{a}/m_{b}=2$, and the momentum cutoff is $\Lambda/k_{F}=10$.
We numerically map out the in-vacuum phase diagram in the plane of $1/(\Lambda a_{s})$ and $1/(\Lambda a_{p})$ as shown in Fig. 3. The in-vacuum phase diagram gives the lowest-lying phase among all the five configurations of two- or three-body clusters discussed in this work. The mass ratio is taken as $m_{a}=2m_{b}$. Within the explored ranges of $a_{s}$ and $a_{p}$ shown in Fig. 3, the in-vacuum three-body states are always more deeply bound than the two-body ones. Here the binding energy is defined as the energy measured relative to the vacuum threshold, so that a more strongly bound state corresponds to a more negative energy (or a larger $|E|$). Physically, the predominance of three-body binding in vacuum can be understood as a cooperative effect of the coexistent $s$- and $p$-wave attractions. While the $s$-wave interaction favors interspecies pairing and the $p$-wave interaction provides attraction between identical fermions, a three-body configuration can benefit from both channels simultaneously. This additional binding mechanism is absent in isolated two-body states, leading to a deeper trimer binding within the explored parameter regime. Moreover, at a fixed $s$-wave interaction, with the increase of $p$-wave interaction, the lowest-lying phase changes from the three-body states with aabaab configuration to those with abbabb configuration. In other words, the in-vacuum three-body states with abbabb configuration become dominant in the regime with stronger $p$-wave interaction. It is because the increase of $p$-wave attraction induces larger enhancement for $bb$ pairings with smaller mass compared to $aa$ ones.